OpenStudy (anonymous):

I don't know how to approach this question (help would be greatly appreciated!) Find the average rate of change of the area of the circle with respect to its radius r as r changes from: (i) 2 to 3, (ii) 2 to 2.5, (iii) 2 to 2.1

OpenStudy (anonymous):

Oh actually I think I got it... I differentiate pi*r^2 and then plug in the two values given and then find the average , right?

OpenStudy (freckles):

\[\frac{\pi \cdot r_1^2 -\pi \cdot r_2^2}{r_1-r_2} \\ \pi \frac{r_1^2-r_2^2}{r_1-r_2} \\ \pi \frac{(r_1-r_2)(r_1+r_2)}{r_1-r_2} \\ \pi (r_1+r_2)\]

OpenStudy (anonymous):

2pi(2) + 2pi(3) ------------- is this right? 2

OpenStudy (anonymous):

for part (i)

OpenStudy (freckles):

\[\frac{\pi \cdot r_1^2 -\pi \cdot r_2^2}{r_1-r_2} \\ \pi \frac{r_1^2-r_2^2}{r_1-r_2} \\ \pi \frac{(r_1-r_2)(r_1+r_2)}{r_1-r_2} \\ \pi (r_1+r_2)\] this is the average rate of change of the area with respect to radius

OpenStudy (freckles):

all you have to do is add your r's and then multiply the result by pi

OpenStudy (anonymous):

Ahhh I see! That makes sense!

OpenStudy (freckles):

like I started with the one formula change of area divided by change or radi

OpenStudy (freckles):

\[\text{ average rate iof change of area}=\frac{A(r_1)-A(r_2)}{r_1-r_2}\]

OpenStudy (freckles):

but was able to simplify this to that really pretty formula

OpenStudy (anonymous):

Right, and that gives us the avg rate of change of area

OpenStudy (anonymous):

Oh that sure is a beautiful formula *_*

OpenStudy (freckles):

right if you wanted instantaneous rate of change at some r you would find the derivative then plug in that r

OpenStudy (anonymous):

Ohhh true!!

OpenStudy (sepeario):

how do you type text like that in the formula thing?

OpenStudy (anonymous):

@freckles thanks again!! :)

OpenStudy (anonymous):

@Sepeario click on 'equation'

OpenStudy (sepeario):

yeah but it always italicises the text

OpenStudy (freckles):

\[\frac{ \text{ kitty eats kitten food} }{\text{ how to do } 5+\frac{6}{7} \text{ in a fraction }}\]

OpenStudy (freckles):

\frac{ \text{ kitty eats kitten food} }{\text{ how to do } 5+\frac{6}{7} \text{ in a fraction }}

OpenStudy (sepeario):

how do you write normal text like that inside the equation thing?

OpenStudy (anonymous):

Ohh hmm I have no clue, @freckles should know :P

OpenStudy (freckles):

:) yes lol  put around that just wherever you want text do \text{ }

OpenStudy (anonymous):

Oooooh how did you know that?? :O

OpenStudy (sepeario):

oh thanks so much

OpenStudy (freckles):

just a little experience with LaTeX

OpenStudy (freckles):

I had to look up some things and also there is a LaTeX group if you want to get really good at it

OpenStudy (anonymous):

Awesomeness ;D

OpenStudy (anonymous):

Oh is there?

OpenStudy (freckles):

http://openstudy.com/study#/groups/LaTeX%20Practicing!%20%3A) there are some tutorials you can find somewhere in these threads they have posted

OpenStudy (anonymous):

Thats great!! Thanks for the link! It'll be useful! :D

OpenStudy (freckles):

By the way I do not claim to be a genius at it

OpenStudy (freckles):

I just know some basic things

OpenStudy (anonymous):

It's still a lot more that I know :P

OpenStudy (anonymous):

than*

OpenStudy (yrelhan4):

The "beautiful formula" thing kinda made me laugh. :P Nice explanation though. :)

OpenStudy (anonymous):

Haha @yrelhan4 yeah same here xD and yes, @freckles explains perfectly!

OpenStudy (freckles):

awww how kind

OpenStudy (anonymous):

^_^ just being honest :D

OpenStudy (freckles):

goodnight and thanks good luck with calculus

OpenStudy (anonymous):

@freckles G'night and thank YOU for helping me out so much! Thanks! :)

OpenStudy (freckles):

It was no problem

OpenStudy (anonymous):

:) Sleep well